# Random Forests as Statistical Procedures: Design, Variance, and Dependence

### v2

@@ Abstract
Random forests are widely used prediction procedures, yet are typically described algorithmically rather than as statistical designs acting on a fixed set of covariates. We develop a finite‑sample, design‑based formulation of random forests in which each tree is an explicit randomized conditional regression function. This perspective yields an exact variance identity for the forest predictor that separates finite‑aggregation variability from a structural dependence term that persists even under infinite aggregation. We further decompose both single‑tree dispersion and inter‑tree covariance using the laws of total variance and covariance, isolating two fundamental design mechanisms—reuse of training observations and alignment of data‑adaptive partitions. These mechanisms induce a strict covariance floor, demonstrating that predictive variability cannot be eliminated by increasing the number of trees alone. The resulting framework clarifies how resampling, feature‑level randomization, and split selection govern resolution, tree variability, and dependence, and establishes random forests as explicit finite‑sample statistical designs whose behavior is determined by their underlying randomized construction.
Keywords: Random Forest; Randomized Design; Variance; Covariance; Resolution
@@ 1.1 Existing theory and its limitations
Since their introduction by Breiman (2001), Random Forests have been among the most widely used modeling frameworks for prediction, adopted across a wide range of scientific domains, and widely regarded as providing strong predictive performance, stability, and robustness to tuning (Hastie et al., 2009). Despite this widespread use, random forests are most often presented as algorithms—ensembles of randomized decision trees motivated empirically or through large-sample arguments—rather than as finite-sample statistical procedures defined by an explicit randomization scheme.
@@
This algorithmic framing has shaped much of the theoretical literature. Early work studied their consistency properties (Breiman, 2004), while subsequent work established consistency and rates of convergence under structural assumptions on splitting rules and feature distributions (Biau et al., 2008; Biau, 2012; Scornet et al., 2015; Biau and Scornet, 2016). Complementary work showed that random forest predictors can be represented as adaptive nearest-neighbor or weighted regression estimators, with predictions expressed as weighted averages of observed responses and weights induced by recursive partitioning (Lin and Jeon, 2006; Biau and Devroye, 2010; Meinshausen, 2006; Athey et al., 2019). These results reveal an underlying statistical structure, but do not provide a finite-sample characterization of how predictive variability and dependence arise from the forest-generating randomization itself.
@@
More recent work has treated random forests as statistical estimators and developed asymptotic distributional theory using U-statistic, V-statistic, and influence-function representations (Mentch and Hooker, 2016; Zhou et al., 2021; Wager and Athey, 2018). These approaches characterize sampling variability under repeated data generation, but they do not analyze the procedural variability induced by the randomization used to construct the forest for a fixed observed dataset. More generally, the variability of a random forest predictor reflects two distinct sources of randomness: sampling variability arising from repeated draws of the dataset, and procedural variability induced by the ‘randomness’ of the Random Forest itself. Existing theory primarily addresses the former. Related work on variance estimation (Xu et al., 2024) conditions on the forest construction mechanism and yields asymptotically valid estimators under repeated sampling, but likewise does not provide an exact finite-sample decomposition of the variability induced by the tree-generating design.
@@
At the same time, it has long been recognized that correlation between ensemble members limits the extent to which aggregation can reduce predictive variability. This phenomenon appears prominently in the literature on bagging and ensemble learning (Breiman, 1996; Dietterich, 2000; Bühlmann and Yu, 2002; Kuncheva and Whitaker, 2003; Brown et al., 2005). More recent work emphasizes that additional sources of randomization can improve predictive performance by weakening dependence between trees, even when individual trees become more variable (Mentch and Zhou, 2020, 2022; Liu and Mazumder, 2025). Other analyses study algorithmic variance and convergence of randomized ensembles conditional on the training data (Lopes et al., 2019a, b). These contributions document the presence and consequences of dependence, but they do not yield an explicit finite-sample decomposition of predictive variability or isolate the structural sources of dependence induced by distinct design choices.
@@ 1.2 A design-based perspective
Recent work has synthesized Random Forests as adaptive smoothing procedures (Curth et al., 2024). We adopt a complementary perspective. We treat a random forest as a finite-sample statistical procedure generated by an explicit randomized design acting on a realized covariate configuration. Specifically, we fix the observed covariate points $X$ and analyze predictive variability under the joint randomness of the outcome realization $(Y\mid X)$ and the tree-generating mechanism $\theta$.
@@
This conditioning choice is deliberate. Fixing $X$ but not $Y$ preserves both sources of variability that affect the forest predictor at the moment it is constructed: the intrinsic dispersion of outcomes at the realized design points, and the algorithmic randomness introduced by the forest construction. A practitioner observes covariate profiles that are determined by the study design or sampling frame; what remains uncertain is both what outcomes those covariate profiles produce and how the tree-building algorithm partitions them. Conditioning on $X$ captures this joint uncertainty.
@@
Within this framework, each tree is viewed as arising from randomized operations—such as observation selection (Breiman, 1996), random subspace selection for candidate splits (Ho, 1998), and stochastic split choice—that determine how the outcomes are locally averaged to form predictions. The tree-generating mechanism induces a probability distribution over terminal regions and weight vectors conditional on $X$, and all algorithmic variability is defined with respect to this design. Outcome randomness enters through the conventional fixed-design model $Y\mid X$. All results are finite-sample and hold at fixed $n$.
@@
When the dataset itself is viewed as random under repeated sampling, the unconditional variability of the forest predictor decomposes into a sampling component and a design component. The sampling component reflects instability of the induced infinite-forest target (the deterministic limit of the forest predictor as the number of trees goes to infinity) across datasets and, under standard stability conditions, decreases at the $O(n^{-1})$ rate. By contrast, the design component captures variability introduced by the forest’s randomized construction acting on the realized covariate configuration and depends explicitly on aggregation, subsampling, and partitioning choices. Section 2 formalizes this decomposition and motivates our focus on the design-induced variability that distinguishes random forests as statistical procedures.
@@
Within this framework, we study the random forest predictor itself as a statistical procedure induced by a randomized design. Although resampling-based performance estimates—most notably out-of-bag evaluation—are produced as a byproduct of the forest construction, they correspond to distinct estimands and should not be conflated with the predictor itself. Out-of-bag evaluation is a form of internal cross-validation targeting predictive risk, whereas our analysis concerns the behavior of the fitted forest predictor itself at a fixed prediction point. This distinction parallels classical regression analysis, in which one studies the properties of a fitted generalized linear model independently of how cross-validation is used to assess its predictive performance.
@@
We show that random forests can be represented as averages of randomized conditional regression functions. Each tree induces a data-adaptive local regression, expressible as a weighted average of observed responses with weights determined by the induced partition of the feature space. Averaging these randomized regressions yields a natural design-based target for the forest predictor. Within this formulation, we derive an exact finite-sample variance identity that separates aggregation variability from a structural dependence component that persists under infinite aggregation, and we show how common forest hyperparameters govern resolution, single-tree variability, and structural dependence.
@@ 2.1 Variance decomposition and scope
Let $\mathcal{D}_{n}^{\star}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ denote a generic sample of size $n$ from the population, and let $\hat{f}_{B}(x;\mathcal{D},\theta)$ denote the $B$-tree forest predictor constructed from dataset $\mathcal{D}$ using design randomization $\theta$. To isolate variability conditional on the realized covariate configuration, we apply the law of total variance with respect to $X$. This yields
@@
The first term represents the finite-sample variability of a forest at a realized covariate configuration, where randomness arises jointly from outcome variation under the fixed-design model $Y\mid X$ and from the tree-generating mechanism $\theta$. The second term reflects sampling variability of the covariate configuration itself; it captures how the induced infinite-forest target changes across repeated draws of $X$ from the population. Under standard stability conditions, this sampling component is generically $O(n^{-1})$; see Appendix A and related results in Mentch and Hooker (2016); Wager and Athey (2018).
@@ 2.2 Tree-level randomized regression functions
We formalize a single tree as a randomized prediction rule acting on a realized covariate configuration $X=\{X_{i}\}_{i=1}^{n}$, with $X_{i}\in\mathcal{X}\subset\mathbb{R}^{p}$. Outcomes satisfy a fixed-design model $Y\mid X$, and randomness in the tree arises jointly from the outcome realization under this model and from the algorithmic construction. A tree is generated by applying a randomized recursive partitioning algorithm to the training sample $(X,Y)$, incorporating operations such as observation selection, random subspace selection for candidate splits, and stochastic split choice. A particular realization of this algorithmic randomization is indexed by $\theta$, and the resulting predictor is denoted by $T_{\theta}$. At a fixed realization $(Y,\theta)$ and fixed covariate configuration $X$, the tree and its induced prediction rule are deterministic functions.
@@
consisting of the training observations grouped with $x$ in the terminal node reached by $x$. The set $A_{\theta}(x)$ is the primitive object governing the tree-level prediction at $x$ and depends on all sources of algorithmic randomness. Define the terminal membership indicator
@@
The weights $W_{i}(x;\theta)$ are normalized membership indicators derived entirely from the averaging set $A_{\theta}(x)$ and introduce no additional randomness beyond that induced by the tree-generating mechanism. As $\theta$ varies under the tree-generating mechanism, the prediction $T_{\theta}(x)$ varies solely through the induced averaging set, equivalently through the collection of indicators $\{M_{i,\theta}(x)\}_{i=1}^{n}$. Although the covariate configuration $X$ is fixed, the averaging set $A_{\theta}(x)$ depends on the realized outcomes $Y$ because split selection optimizes impurity criteria computed from the training responses available to tree $\theta$.
Representations of random forests as adaptive weighted averages of the training responses are well established (Lin and Jeon, 2006; Biau and Devroye, 2010; Scornet et al., 2015) and underlie extensions such as quantile regression forests (Meinshausen, 2006). Our contribution is not the algebraic form itself, but the probabilistic framing: we treat the random weight vector $W(x;\theta)$ as the primary random object in the induced probability space. In this view, the infinite-aggregation predictor arises as the expectation of this random weight vector under the tree-generating mechanism conditional on $X$, and subsequent variance properties are governed by its distribution.
@@ 2.3 The forest predictor and its design-based target
A random forest predictor is a Monte Carlo average of randomized tree-level prediction rules. Let $\{T_{\theta_{b}}\}_{b=1}^{B}$ denote $B$ independent realizations of the tree-generating mechanism described in Section 2.2. The finite forest predictor is
@@
The aggregation level $B$ controls only the amount of Monte Carlo averaging and does not alter the distribution of the underlying tree-level rule. Let $\theta$ denote a generic draw from the tree-generating mechanism and define the infinite-aggregation target as the $\theta$-average conditional on the realized covariate configuration,
@@
Because $X$ is fixed but the training outcomes remain random under the fixed-design model, $f_{\infty}(x)$ is random through $Y$. The finite forest predictor $\hat{f}_{B}(x)$ is a Monte Carlo approximation to $f_{\infty}(x)$ and satisfies $\mathbb{E}[\hat{f}_{B}(x)\mid X]=f_{\infty}(x)$. Finite forests fluctuate around this target due to limited aggregation, while structural dependence between tree-level prediction rules induces a non-vanishing covariance component as $B\to\infty$. The variance analysis that follows makes this separation explicit.
@@ 3.1 Finite-sample variance identity
We now characterize the finite-sample variance of the random forest predictor $\hat{f}_{B}(x)$ conditional on the realized covariate configuration $X$. Recall that $\hat{f}_{B}(x)$ is the average of $B$ independent realizations $T_{\theta_{b}}(x)$ of the randomized tree-level prediction rule $T_{\theta}(x)$, where randomness arises jointly from $(Y\mid X)$ and the tree-generating mechanism. Define
@@
where $\theta$ and $\theta^{\prime}$ denote independent draws from the tree-generating mechanism. Although trees are generated independently, the resulting predictions need not be independent, because each tree acts on the same realized dataset through design-induced random averaging sets. This dependence is captured by the covariance term $C_{T}(x)$.
@@
This identity separates variability due to finite aggregation, represented by $\sigma_{T}^{2}(x)/B$, from structural dependence induced by the design through the joint behavior of tree-level averaging sets, represented by $C_{T}(x)$. The result follows directly from exchangeability of the tree-level prediction rules under the tree-generating mechanism.
@@ 3.2 Decomposing the single-tree variance
We now decompose the single-tree variance term $\sigma_{T}^{2}(x)=\operatorname{Var}(T_{\theta}(x)\mid X)$ appearing in Theorem 1. Let $\mathcal{I}_{\theta}=(I_{\theta,1},\dots,I_{\theta,n})$ denote the resampling indicators for the tree indexed by $\theta$, where $I_{\theta,i}=1$ if observation $i$ is exposed to the tree construction and $I_{\theta,i}=0$ otherwise. For a fixed prediction point $x$, the resampling indicators determine which observations are eligible to appear in the random averaging set $A_{\theta}(x)$, while the remaining algorithmic randomness—such as feature subsampling and split selection—determines how $x$ is grouped with those observations. Conditioning on $\mathcal{I}_{\theta}$ in addition to $X$ and applying the law of total variance yields
$\operatorname{Var}\!\bigl(T_{\theta}(x)\mid X\bigr)=\mathbb{E}\!\left[\operatorname{Var}\!\bigl(T_{\theta}(x)\mid\mathcal{I}_{\theta},X\bigr)\mid X\right]+\operatorname{Var}\!\left(\mathbb{E}\!\bigl[T_{\theta}(x)\mid\mathcal{I}_{\theta},X\bigr]\mid X\right).$ (1)
The first term in (1) captures variability arising from random splitting decisions after the resampling pattern is fixed, while the second captures variability induced by changes in the resampling realization itself, which alter which observations are eligible to be grouped with $x$. Both components are evaluated under the joint randomness of $(Y\mid X)$ and the tree-generating mechanism $\theta$, with $X$ treated as fixed throughout. We now focus on the first component and examine how randomness in local grouping contributes to single-tree variability.

### v3

# arXiv:2602.13104v3[stat.ML] 28 Feb 2026

@@ Abstract
We develop a finite-sample, design-based theory for random forests in which each tree is a randomized conditional predictor acting on fixed covariates and the forest is their Monte Carlo average. An exact variance identity separates Monte Carlo error from a covariance floor that persists under infinite aggregation. The floor arises through two mechanisms: observation reuse, where the same training outcomes receive weight across multiple trees, and partition alignment, where independently generated trees discover similar conditional prediction rules. We prove the floor is strictly positive under minimal conditions and show that alignment persists even when sample splitting eliminates observation overlap entirely. We introduce procedure-aligned synthetic resampling (PASR) to estimate the covariance floor, decomposing the total prediction uncertainty of a deployed forest into interpretable components. For continuous outcomes, resulting prediction intervals achieve nominal coverage with a theoretically guaranteed conservative bias direction. For classification forests, the PASR estimator is asymptotically unbiased, providing the first pointwise confidence intervals for predicted conditional probabilities from a deployed forest. Nominal coverage is maintained across a range of design configurations for both outcome types, including high-dimensional settings. The underlying theory extends to any tree-based ensemble with an exchangeable tree-generating mechanism.
Keywords: Random Forest; Variance Decomposition; Covariance Floor; Design-Based Inference; Prediction Intervals; Tree Ensembles
@@ 1.1 Existing theory and its limitations
Since their introduction by Breiman (2001), Random Forests have been among the most widely used modeling frameworks for prediction, adopted across a wide range of scientific domains, and widely regarded for providing strong predictive performance, stability, and robustness to tuning (Hastie et al., 2009). Despite their ubiquity in applied statistical settings, Random Forests are typically presented as algorithms—ensembles of randomized decision trees motivated empirically or through large-sample arguments—rather than as finite-sample statistical procedures.
@@
This algorithmic framing has shaped much of the theoretical literature. Early work studied consistency properties (Breiman, 2004), and subsequent results established consistency and rates of convergence under structural assumptions on splitting rules and feature distributions (Biau et al., 2008; Biau, 2012; Scornet et al., 2015; Biau and Scornet, 2016). Complementary work showed that random forest predictors can be represented as adaptive nearest-neighbor or weighted regression estimators, with predictions expressed as weighted averages of observed responses and weights induced by recursive partitioning (Lin and Jeon, 2006; Biau and Devroye, 2010; Meinshausen, 2006; Athey et al., 2019). More recent work has treated random forests as statistical estimators and developed asymptotic distributional theory using U-statistic, V-statistic, and influence-function representations (Mentch and Hooker, 2016; Zhou et al., 2021; Wager and Athey, 2018). Related work on variance estimation (Xu et al., 2024) conditions on the forest construction mechanism and yields asymptotically valid estimators under repeated sampling.
@@
These approaches characterize sampling variability—how the forest predictor would change under repeated draws of the training data from the population—but they do not analyze the procedural variability induced by the randomization used to construct the forest at a fixed observed dataset. More generally, the variability of a random forest predictor reflects two distinct sources of randomness: sampling variability arising from repeated draws of the dataset, and procedural variability induced by the forest’s own randomized construction. Existing theory primarily addresses the former.
@@
It has also long been recognized that correlation between ensemble members limits the extent to which aggregation can reduce predictive variability (Breiman, 1996; Dietterich, 2000; Bühlmann and Yu, 2002; Kuncheva and Whitaker, 2003; Brown et al., 2005). More recent work emphasizes that additional sources of randomization can improve predictive performance by weakening dependence between trees, even when individual trees become more variable (Mentch and Zhou, 2020, 2022; Liu and Mazumder, 2025), and other analyses study algorithmic variance and convergence of randomized ensembles conditional on the training data (Lopes et al., 2020, 2019). These contributions document the presence and consequences of dependence, but do not yield an explicit finite-sample decomposition of predictive variability or isolate the structural sources of dependence induced by distinct design choices.
As a practical consequence, no existing method provides pointwise variance estimates for a deployed random forest conditional on the data used to construct it. Existing asymptotic approaches—including the infinitesimal jackknife (Wager et al., 2014), U-statistic (Mentch and Hooker, 2016), and V-statistic (Zhou et al., 2021) frameworks—are developed for regression-type targets and quantify sampling variability of the infinite-aggregation target across repeated datasets, answering how the forest’s target would change if new training data were collected. Conformal prediction methods (Lei et al., 2018; Romano et al., 2019) provide distribution-free coverage guarantees but do not decompose the sources of predictive variability or characterize the forest’s own design-induced uncertainty. For a practitioner with a fitted forest in hand, none of these approaches quantify the total uncertainty of the delivered prediction at a given covariate value. For classification tasks that yield predicted probabilities, the asymptotic regression machinery can be applied with binary outcomes encoded as $Y\in\{0,1\}$; however, existing theory is formulated for regression-type targets and does not provide pointwise uncertainty quantification for the estimated conditional probability $\hat{p}(x)$ from a deployed forest.
@@ 1.2 A design-based perspective
Recent work has synthesized Random Forests as adaptive smoothing procedures (Curth et al., 2024). We adopt a complementary perspective: we treat a random forest as a finite-sample statistical procedure generated by an explicit randomized design acting on a realized covariate configuration. Specifically, we fix the observed covariate points $X$ and analyze predictive variability under the joint randomness of the outcome realization $(Y\mid X)$ and the tree-generating mechanism $\theta$.
@@
This conditioning choice is deliberate. Fixing $X$ but not $Y$ preserves both sources of variability that affect the forest predictor at the moment it is constructed: the intrinsic dispersion of outcomes at realized covariates $X$, and algorithmic randomness introduced by the forest procedure. A practitioner observes a fixed set of covariate profiles determined by the study design; what remains uncertain is both what outcomes those profiles produce and how the tree-building algorithm partitions them. Conditioning on $X$ captures this joint uncertainty.
@@
Within this framework, each tree arises from randomized operations—observation selection (Breiman, 1996), random subspace selection for candidate splits (Ho, 1998), and random split choice—that determine how outcomes are locally averaged to form predictions. The tree-generating mechanism induces a probability distribution over terminal regions and local averaging rules conditional on $X$, and all algorithmic variability is defined with respect to this design. Outcome randomness enters through the conventional fixed-design model $Y\mid X$. All results are finite-sample and hold at fixed $n$.
@@
When the dataset itself is viewed as random under repeated sampling, the unconditional variability of the forest predictor decomposes into a sampling component and a design component. The sampling component reflects instability of the induced infinite-forest target across datasets and, under standard stability conditions, decreases at the $O(n^{-1})$ rate. The design component captures variability introduced by the forest’s randomized construction acting on the realized covariate configuration and depends explicitly on aggregation, subsampling, and partitioning choices. Section 2 formalizes this decomposition and motivates our focus on the design-induced variability that distinguishes random forests as statistical procedures.
@@
We study the random forest predictor itself, not resampling-based performance estimates such as out-of-bag evaluation. Out-of-bag evaluation is a form of internal cross-validation targeting predictive risk; our analysis concerns the behavior of the fitted predictor at a fixed prediction point. This parallels classical regression analysis, in which one studies properties of a fitted model independently of how cross-validation is used to assess its predictive performance.
@@
We show that each tree induces a data-adaptive conditional regression, which can be written as as a weighted average of observed responses, and that the forest predictor is an average of these randomized regression functions. This formulation yields an exact finite-sample variance identity that separates aggregation variability from a structural dependence component that persists under infinite aggregation, and reveals how common forest hyperparameters govern resolution, single-tree variability, and structural dependence.
@@ 2.1 Variance decomposition and scope
Let $\mathcal{D}_{n}^{\star}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ denote a generic sample of size $n$ from the population, and let $\hat{f}_{B}(x;\mathcal{D},\theta)$ denote the $B$-tree forest predictor constructed from dataset $\mathcal{D}$ using design randomization $\theta$. To isolate variability conditional on the realized covariate configuration, we apply the law of total variance with respect to $X$:
@@
The first term represents the finite-sample variability of a forest at a realized covariate configuration, where randomness arises jointly from outcome variation under the fixed-design model $Y\mid X$ and from the tree-generating mechanism $\theta$. The second term reflects sampling variability of the covariate configuration itself; it captures how the induced infinite-forest target changes across repeated draws of $X$ from the population. Under standard stability conditions, this sampling component is generically $O(n^{-1})$; see Appendix A.1 and related results in Mentch and Hooker (2016); Wager and Athey (2018).
@@ 2.2 Tree-level randomized regression functions
We formalize a single tree as a randomized conditional predictor acting on a realized covariate configuration $X=\{X_{i}\}_{i=1}^{n}$, with $X_{i}\in\mathcal{X}\subset\mathbb{R}^{p}$. Outcomes satisfy a fixed-design model $Y\mid X$, and randomness in the tree arises jointly from the outcome realization and from the algorithmic construction. A particular realization of this randomization is indexed by $\theta$, and the resulting predictor is denoted by $T_{\theta}$. At a fixed realization $(Y,\theta)$ and covariate configuration $X$, the tree and its induced conditional predictor are deterministic functions.
@@
consisting of the training observations grouped with $x$ in the terminal node reached by $x$. The set $A_{\theta}(x)$ is the primitive object governing the tree-level prediction at $x$ and depends on all sources of algorithmic randomness. Define the terminal node membership indicator
@@
The weights $W_{i}(x;\theta)$ are normalized membership indicators derived entirely from the averaging set $A_{\theta}(x)$. As $\theta$ varies under the tree-generating mechanism, the prediction $T_{\theta}(x)$ varies solely through the induced averaging set, equivalently through the collection of indicators $\{M_{i,\theta}(x)\}_{i=1}^{n}$. Although the covariate configuration $X$ is fixed, the averaging set $A_{\theta}(x)$ depends on the realized outcomes $Y$ because split selection optimizes impurity criteria computed from the training responses available to tree $\theta$.
Representations of random forests as adaptive weighted averages of the training responses are well established (Lin and Jeon, 2006; Biau and Devroye, 2010; Scornet et al., 2015) and underlie extensions such as quantile regression forests (Meinshausen, 2006). Our contribution is not this algebraic form, but the probabilistic framing: we treat the random weight vector $W(x;\theta)$ as the primary random variable of interest. The infinite-aggregation predictor is the expectation of this random weight vector under the tree-generating mechanism conditional on $X$, and subsequent variance properties are determined by its distribution.
@@ 2.3 The forest predictor and its design-based target
A random forest predictor is a Monte Carlo average of randomized tree-level conditional predictors. Let $\{T_{\theta_{b}}\}_{b=1}^{B}$ denote $B$ independent realizations of the tree-generating mechanism described in Section 2.2. The finite forest predictor is
@@
The aggregation level $B$ controls only the amount of Monte Carlo averaging and does not alter the distribution of the underlying tree-level predictor. Let $\theta$ denote a generic draw from the tree-generating mechanism and define the infinite-aggregation target as the $\theta$-average conditional on the realized covariate configuration,
@@
Because $X$ is fixed but the training outcomes remain random under the fixed-design model, $f_{\infty}(x)$ is random through $Y$. The finite forest predictor $\hat{f}_{B}(x)$ is a Monte Carlo approximation to $f_{\infty}(x)$ and satisfies $\mathbb{E}[\hat{f}_{B}(x)\mid X]=f_{\infty}(x)$. Finite forests fluctuate around this target due to limited aggregation, while structural dependence between tree-level conditional predictors induces a non-zero covariance component as $B\to\infty$. The variance analysis that follows makes this separation explicit.
@@ 3.1 Finite-sample variance identity
We now characterize the finite-sample variance of the random forest predictor $\hat{f}_{B}(x)$ conditional on the realized covariate configuration $X$. Define
@@
where $\theta$ and $\theta^{\prime}$ denote independent draws from the tree-generating mechanism. Although trees are generated independently, the resulting predictions are not independent, because each tree averages outcomes from the same realized dataset and the same observations can contribute to predictions across multiple trees. This dependence is captured by the covariance term $C_{T}(x)$.
@@
This identity separates variability due to finite aggregation, $\sigma_{T}^{2}(x)/B$, from structural dependence induced by the design, $C_{T}(x)$.
@@ 3.2 Decomposing the single-tree variance
We now decompose the single-tree variance term $\sigma_{T}^{2}(x)=\operatorname{Var}(T_{\theta}(x)\mid X)$ appearing in Theorem 1. Let $\mathcal{I}_{\theta}=(I_{\theta,1},\dots,I_{\theta,n})$ denote the resampling indicators for the tree indexed by $\theta$, where $I_{\theta,i}=1$ if observation $i$ is exposed to the tree construction and $I_{\theta,i}=0$ otherwise. The resampling indicators determine which observations are eligible to appear in the random averaging set $A_{\theta}(x)$, while the remaining algorithmic randomness—feature subsampling and split selection—determines how $x$ is grouped with those observations. We condition on $\mathcal{I}_{\theta}$ in addition to $X$ to separate these two sources of variability; this follows the operational structure of the algorithm, in which the resampling step determines which observations are available to each tree before any splitting decisions are made. Applying the law of total variance yields
$\operatorname{Var}\!\bigl(T_{\theta}(x)\mid X\bigr)=\underbrace{\mathbb{E}\!\left[\operatorname{Var}\!\bigl(T_{\theta}(x)\mid\mathcal{I}_{\theta},X\bigr)\mid X\right]}_{V_{\mathrm{in}}(x)}+\underbrace{\operatorname{Var}\!\left(\mathbb{E}\!\bigl[T_{\theta}(x)\mid\mathcal{I}_{\theta},X\bigr]\mid X\right)}_{V_{\mathrm{out}}(x)}.$ (1)
The first term captures variability arising from random splitting decisions after the resampling pattern is fixed; the second captures variability induced by changes in the resampling realization itself. Both components are evaluated under the joint randomness of $(Y\mid X)$ and the tree-generating mechanism $\theta$.